OpenStudy (anonymous):

6x = 7 - 4y; x= -2,-1, 0 SHOW WORK PLEASE

OpenStudy (anonymous):

x=-2 6(-2)=7-4y -12=7-4y -19=-4y 19/4=y x=-1 6(-1)=7-4y -6=7-4y -13=-4y 13/4=y x=0 6(0)=7-4y 0=7-4y -7=-4y 7/4=y